If the sum of first n natural numbers is 1/5 times the sum of their squares, then th value of n is

Grade:11

1 Answers

Ajay
7 years ago
This question should have been posted in algebra section . Anyway here is solution.
It was easy, wonder why you cant do it
sum\quad of\quad n\quad natural\quad numbers\quad =\quad \frac { n(n+1) }{ 2 } \\ Sum\quad of\quad squares\quad of\quad n\quad natural\quad numbers\quad =\quad \frac { n(n+1)(2n+1) }{ 6 } \\ Given\quad \frac { n(n+1) }{ 2 } \quad =\quad \frac { 1 }{ 5 } \left( \frac { n(n+1)(2n+1) }{ 6 } \right) \\ solving\quad for\quad n,\quad n\quad =7